OpenStudy (anonymous):

Solve the system: X+3y = 1 -2x + y = -23 Answers: x = 16; y = –5 x = 10; y = –3 x = 22; y = 7 x = 14; y = 5

OpenStudy (jjuden):

https://www.mathway.com/

OpenStudy (hockeychick23):

you can use elimination and solve this very easily, you need to get the 1st to have 2x just like the 2nd one. (x+3y=1)2 -2x+y=-23 Now you have 2 similar equations: 2x+6y=2 -2x+y= -23 Add them together 7y= -21 y=-3 Plug -3 into one of the equations x+3(-3)=1 x-9=1 x=10 So x=10, y= -3

OpenStudy (anonymous):

Thank you or actually explaining how to acheive the answer as well. Thanks!!

OpenStudy (hockeychick23):

no problem!(:

OpenStudy (funnyguy2):

stuff